3. A loan of $36,550.00 at 5.00% compounded semi-annually is to be repaid with payments at the end of every 6 months. The loan was settled in 5 years.

 

a. Calculate the size of the periodic payment. Round to the nearest cent.

 b. Calculate the total interest paid. Round to the nearest cent.


1
Expert's answer
2021-08-12T15:02:35-0400

a)

Interest rate =5%

Semi-annual interest rates "=\frac{5}{2}=2.5\%"

Payment period =5 years =10 semi-annual

Present value FACTOR "=\frac{1-(1+r)^{-n}}{r}"  

Present value FACTOR "=\frac{1-(1.025)^{-10}}{0.025}=8.752063931"

Loan amount =Present value FACTOR x semi-annual payment 

"36550=8.7520663931 \u00d7" semi-annual payment 

Semi-annual payment "=\$4176.157"

Total payment = Semi-annual payment x number of payments 

Total payment "=10 \u00d7 4176.157=41761.57"

Total payment "=\$41761.57"

b)

"Interest =Total \space payment- loan \space amount \\

Interest =41761.57 - 36550\\

Interest =\$5211.57\\"